# Optimization under rare events: scaling laws for linear chance-constrained programs

### v1

@@ 7.5. Proof of Theorem 5
The constraint in ($\rm{CVaR}_{\delta}$) with $x=y/r_{\delta}$ is equivalent to
${\bf E}\big{[}\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{]}\leq 1$ (44)
where $\tau_{\delta}$ is the optimizing value in (5). Using (12), it satisfies
$\delta={\bf P}(\phi(y/r_{\delta},L)-1>\tau_{\delta})$ $={\bf P}\big{(}\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{)}$
$=(1+o(1)){\bf E}[\phi(y,\Theta)^{\alpha}]{\bf P}(|L|\geq r_{\delta}(\tau_{\delta}+1)).$
As a result, we obtain that $(\tau_{\delta}+1)r_{\delta}=\bar{F}^{-1}(\frac{(1+o(1))\delta}{{\bf E}[\phi(y,\Theta)^{\alpha}]})\to\infty$ as $\delta\downarrow 0$. We substitute this in the left-hand side in (44), and recall that $\bar{F}^{-1}(\delta)=r_{\delta}$ by definition. Then, as $\delta\downarrow 0$,
${\bf E}\big{[}\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{]}$ $=\frac{\bar{F}^{-1}\Big{(}\frac{(1+o(1))\delta}{{\bf E}[\phi(y,\Theta)^{\alpha}]}\Big{)}}{\bar{F}^{-1}(\delta)}\cdot{\bf E}\left[\frac{\phi(y,L)}{r_{\delta}(\tau_{\delta}+1)}\,\Big{|}\,\frac{\phi(y,L)}{r_{\delta}(\tau_{\delta}+1)}>1\right]$
$\rightarrow({\bf E}[\phi(y,\Theta)^{\alpha}])^{1/\alpha}\cdot\frac{\alpha}{\alpha-1}.$
The convergence of the first factor follows from regular variation. The convergence of the second factor holds because $\phi(y,L)/r_{\delta}(\tau_{\delta}+1)$, conditional upon exceeding 1, converges to a Pareto distributed random variable on $[1,\infty)$ with exponent $\alpha$. This convergence also holds in expectation; to show this we can apply a Potter bound for the ratio of regularly varying functions as we did in the proof of Theorem 2.
We now apply Proposition 1 with the values $k=1/\delta$, $g_{k}(y)={\bf E}[\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)]$, and $g(y)=({\bf E}[\phi(y,\Theta)^{\alpha}])^{1/\alpha}\frac{\alpha}{\alpha-1}$. We conclude that the normalized version of ($\rm{CVaR}_{\delta}$) converges to the optimal value of
maximize $\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq(1-1/\alpha)^{\alpha}.$ (45)
Noting that ($\rm{HT}$) has a unique solution $y^{*}$, we conclude that the optimal solution to (45) equals $y^{*}(1-1/\alpha)$ by the homogeneity of $\phi$. Combining this with Theorem 4 completes the proof. $\square$
@@ 7.6. Proof of Theorem 6
Write $a_{k}:=\bar{F}^{-1}(1/k)$ for the normalizing constants in (${\rm{SCP}\textnormal{-}\rm{HT}}_{k}$). To show convergence in distribution of the optimal value $V_{SCP}(k)$ to $V^{\ast}$, it suffices to show that the non-feasible set spanned by the constraints $\phi(y,L^{(i)}/a_{k})\leq 1$, $i=1,\ldots,k$, converges weakly to the non-feasible set spanned by the constraints $\phi(y,z)\leq 1$, $z\in N$, where $N$ is the Poisson point process defined above (14). To do so, we rely on the theory of random sets, as developed in Molchanov (2017).
@@
where we use the notation $bS=\{x:bx\in S\}$. Applying Theorem 4.2.9 and 4.2.10 of Molchanov (2017), we obtain that $U_{k}$ converges in distribution to $U=\{y:\phi(y,z)\geq 1\mbox{ for some }z\in N\}$. We define for $t>0$ the half-space $K_{t}=\{y:c^{T}y>t\}$. Therefore, we see that for $t>0$,

### v2

@@ 8.5. Proof of Theorem 5
Similar to (28), we write (5) with $x=y/r_{\delta}$ as
${\bf CVaR}(\phi(y/r_{\delta},L))=\inf_{\tau\in\mathbb{R}}\left\{\tau+\frac{1}{\delta}\int_{\tau}^{\infty}{\bf P}(\phi(y,L)>r_{\delta}u)du\right\}=\tau_{\delta}+\frac{1}{r_{\delta}\delta}\int_{r_{\delta}\tau_{\delta}}^{\infty}{\bf P}(\phi(y,L)>v)dv,$ (42)
where $\tau_{\delta}=\inf\{t:{\bf P}(\phi(y,L)>r_{\delta}t)\leq\delta\}$ attains the infimum. By the reasoning in the proof of Theorem 4 from (37) until (41) with $y$ replaced by $y/t$, we obtain
${\bf P}(\phi(y,L)>r_{\delta}t)\sim{\bf E}[\phi(y/t,L)^{\alpha}]\bar{F}(r_{\delta})\sim{\bf E}[\phi(y,L)^{\alpha}]t^{-\alpha}\delta.$ (43)
Consequently, $\tau_{\delta}\sim{\bf E}[\phi(y,L)^{\alpha}]^{1/\alpha}$. Since $r_{\delta}\tau_{\delta}\rightarrow\infty$ and ${\bf P}(\phi(y,L)>v)$ is regularly varying with index $\alpha<-1$ by the homogeneity of $\phi$, we can apply Karamata’s theorem to obtain
${\bf CVaR}(\phi(y/r_{\delta},L))\sim\tau_{\delta}+\frac{r_{\delta}\tau_{\delta}}{r_{\delta}\delta}\frac{1}{\alpha-1}{\bf P}(\phi(y,L)>r_{\delta}\tau_{\delta})\sim\frac{\alpha}{\alpha-1}\tau_{\delta}\sim\frac{\alpha}{\alpha-1}{\bf E}[\phi(y,L)^{\alpha}]^{1/\alpha}.$ (44)
Now, we apply Proposition 2 with $k=1/\delta$, $g_{k}(y)={\bf CVaR}(\phi(y/r_{\delta},L))$, and $g(y)=\frac{\alpha}{\alpha-1}{\bf E}[\phi(y,L)^{\alpha}]^{1/\alpha}$. From the variational representation of ${\bf CVaR}(\phi(y/r_{\delta},L))$ and the fact that $\phi$ is nondecreasing in $y$ for every fixed $L$, it follows that $g_{k}$ is nondecreasing in $y$, and $g(ay)$ is strictly increasing in $a$ due to the scaling property of $\phi$. Since we also have established that $g_{k}(y)\rightarrow g(y)$, we can apply Proposition 1 to conclude that the normalized version of ($\rm{CVaR}_{\delta}$) converges to the optimal value of
$\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq(1-1/\alpha)^{\alpha}.$ (45)
If $y^{\ast}$ solves ($\rm{HT}$), then $(1-1/\alpha)y^{\ast}$ is an optimal solution of (45) by the homogeneity of $\phi$. Combining this with Theorem 4 completes the proof. $\square$
@@ 8.6. Proof of Theorem 6
Write $a_{k}:=\bar{F}^{-1}(1/k)$ for the normalizing constants in (${\rm{SCP}\textnormal{-}\rm{HT}}_{k}$). To show convergence in distribution of the optimal value $V_{SCP}(k)$ to $V^{\ast}$, it suffices to show that the non-feasible set spanned by the constraints $\phi(y,L^{(i)}/a_{k})\leq 1$, $i=1,\ldots,k$, converges in distribution to the non-feasible set spanned by the constraints $\phi(y,z)\leq 1$, $z\in N$, where $N$ is the Poisson point process defined above (14). To do so, we rely on the theory of random sets, as developed in Molchanov (2017).
@@
where we use the notation $bS=\{x:b^{T}x\in S\}$. Applying Theorems 4.2.9 and 4.2.10 of Molchanov (2017), we obtain that $U_{k}$ converges in distribution to $U=\{y:\phi(y,z)\geq 1\mbox{ for some }z\in N\}$. We define for $t>0$ the half-space $K_{t}=\{y:c^{T}y>t\}$. Therefore, we see that for $t>0$,